OpenStudy (anonymous):

Consider the equation 5x + 2y = 30. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. can anyone help with this?

OpenStudy (anonymous):

part 1: change the equation to slope intercept method by first subrtacting 5x from both sides to get 2y=-5x+30. then divide everything by 2 to get y=-5/2x+15. Then plot your y intercept, which is (0,15) and use your slope, -5/2 to continue drawing points on your line by moving 5 units down and 2 units to the right

OpenStudy (anonymous):

part 2: to use the intercepts method to graph this equation we must first find the intercepts of this graph. to do so, first we set x to be 0 and solve the equation for y. we get 5(0)+2y=30 so ur y intercept is 15 and at the point (0,15). then we set y to be 0 and solve for the x intercept. we get 5x+2(0)=30 which gives us an x intercept at (6,0). then graph the two points and draw a line connecting them